OpenStudy (anonymous):

how do I solve for this equation: e^-4x-3e^-3x = 0

OpenStudy (anonymous):

and I also need to graph the answer

OpenStudy (anonymous):

factor

OpenStudy (anonymous):

\[e^{-4x}(1-3e^x)=0\]

OpenStudy (anonymous):

first factor is never zero, so your job is to solve \[1-3e^x=0\] for \(x\)

OpenStudy (anonymous):

i get stuck at \[e^x = 1/3\]